A 96.6lb is fastened to a cord which is wrapped around a solid cylinder of 3 ft radius weighinh 322lb. The cylinder will raise the weight with an acceleration of 12 ft per seconds? What will be the total bearing reaction?



1
Expert's answer
2021-05-19T06:14:03-0400

Newtons Law of Motion for m1

"T=m_1g+m_1a"

Newtons Law of Motion for m2

"N=m_2g+T"

"N=m_2g+m_1g+m_1a"

"N(lb)=\frac{m_2g+m_1g+m_1a}{g}"

"N(lb)=m_2+m_1+\frac{m_1a}{g}"

"N(lb)=322lb+96.6lb+\frac{96.6*12}{32.174}lb"

"N(lb)=454.62g lb =454.62*32.174lbft\/s^2=14627.236 pound-force"